### v2

# AccidentSim: Generating Vehicle Collision Videos with Physically Realistic Collision Trajectories from Real-World Accident Reports

Abstract
Collecting real-world vehicle accident videos for autonomous driving research is challenging due to their rarity and complexity. While existing video generation methods can produce visually realistic content, they often lack physical realism and fail to generate accurate post-collision vehicle trajectories. In this paper, we introduce AccidentSim, a framework designed to generate physically realistic vehicle collision videos. Specifically, AccidentSim extracts and utilizes physical cues and contextual information from real-world accident reports. It then uses a reliable physical simulator to replicate post-collision vehicle trajectories based on the information from these reports, creating a collision trajectory dataset. This dataset is subsequently used to fine-tune a language model, AccidentLLM, enabling it to predict novel, physically consistent post-collision trajectories across various driving scenarios based on user-provided accident descriptions. Finally, we synthesize collision videos by compositing the predicted trajectories onto vehicles with rendered backgrounds. Experimental results demonstrate that videos generated by AccidentSim excel in both visual and physical realism, and incorporating these videos into autonomous system training can effectively reduce collision rates.
Figure 1: AccidentSim generates vehicle collision videos from user-provided accident descriptions, producing physically realistic collision trajectories of vehicles across diverse collision and road types, all according to the user-specified viewpoints. Particularly, AccidentSim extracts physical cues and contextual information from real-world accident reports and feeds them into a physical simulator to generate physically constrained collision trajectories. These trajectories are then used to fine-tune a language model, termed AccidentLLM, enabling it to predict novel, physically consistent trajectories directly from the collision information extracted from the user-provided accident descriptions, extending beyond the collision scenarios contained in the original accident reports.
1 Introduction
Research and development in autonomous driving critically depend on access to diverse and comprehensive training data to ensure robust performance across various driving scenarios. However, existing datasets, such as KITTI [13], Cityscapes [8], nuScenes [5], Waymo [27], and BDD100K [36], primarily capture normal driving conditions, lacking safety-critical events such as vehicle collisions. This data imbalance arises from the inherent difficulty and prohibitive cost of collecting rare but safety-critical scenarios in real-world environments. This phenomenon is known as the Curse of Rarity (CoR) [20], which refers to the low occurrence rate of safety-critical events in high-dimensional spaces. In autonomous driving systems, CoR significantly limits the ability of deep learning models to learn effective responses to these rare but crucial situations, ultimately undermining the overall reliability and safety of such systems.
With recent advances in video generation models, synthesizing vehicle collision videos has become a feasible, low-cost, and promising alternative to expensive real-world data collection. Such generated videos, if physically realistic, could augment existing datasets and enhance the training of autonomous driving systems, ultimately improving reliability and safety in critical scenarios. However, current video generation methods, including state-of-the-art models such as Kling [18], Gen-4 [25], and Jimeng [4], as well as specialized driving-scenario generators like DriveDreamer [30], Panacea [32], Cosmos [2], and HoloDrive [33], still struggle to produce physically realistic vehicle collisions. In particular, these approaches often fail to reproduce physically realistic collision trajectories that conform to real-world physical constraints and dynamics. This limitation arises because these video generation methods are not explicitly designed to incorporate essential physical cues and constraints. For instance, existing models typically overlook key physical factors, such as vehicle momentum, impulse, and mass distribution, that are critical for accurately modeling real-world collision behaviors. Consequently, while the videos generated by these models may be visually realistic, they lack physical realism, failing to capture the true underlying dynamics of vehicle collisions.
To address these challenges, we propose AccidentSim, a framework that introduces a novel approach to generating physically realistic vehicle collision videos based on real-world accident reports. For instance, accident reports from the National Motor Vehicle Crash Causation Survey (NMVCCS) [1] in the United States, document actual vehicle collisions across diverse vehicle models, collision types, road types, and traffic conditions. AccidentSim leverages the rich physical cues and contextual information embedded in such reports to synthesize vehicle-collision videos with physically realistic post-collision trajectories. As illustrated in Figure 1, AccidentSim first extracts key physical cues, such as vehicle speed, collision angle, and mass, together with contextual information, including road type and collision type, from the accident reports. These cues serve as the physical foundation for constructing realistic collision simulations. The extracted physical cues are then fed into a physical simulator, such as CARLA [10], which enforces corresponding physical constraints to ensure that the simulated post-collision vehicle trajectories adhere to real-world dynamics. However, physical simulators have inherent limitations. For example, physical simulators are typically computationally expensive and require fully specified physical parameters, making them unable to directly respond to user-provided accident descriptions for generating diverse post-collision trajectories. To overcome these challenges, we first construct a dataset from the collision trajectories generated by the simulator using real-world accident reports. This dataset is then used to fine-tune a large language model, termed AccidentLLM, within the AccidentSim framework. As a result, AccidentLLM can generate novel, physically realistic collision trajectories directly from the collision information extracted from the user-provided accident descriptions, extending beyond the scenarios captured in existing accident reports. In this way, AccidentLLM effectively bypasses the expensive simulation process through physics distillation, learning to translate under-specified semantic inputs into physically realistic trajectories efficiently and realistically. Finally, AccidentSim synthesizes collision videos by compositing the predicted trajectories onto vehicles with rendered backgrounds, according to user-specified viewpoints.
In summary, AccidentSim integrates physically grounded cues with physical constraints to generate videos that faithfully replicate diverse real-world vehicle collisions. By bridging the gap between synthetic and real-world scenarios, AccidentSim offers a scalable solution to the scarcity of physically realistic collision data for training and evaluating autonomous driving systems.
Our contributions are three-fold:
  • We propose AccidentSim, a framework that extracts and leverages physical cues and contextual information from real-world accident reports, to generate vehicle collision videos with physical simulation, ensuring physically realistic post-collision trajectories.
  • We develop AccidentLLM, a language-driven module fine-tuned on simulated collision trajectories for pre-collision planning and post-collision prediction, enabling the generation of novel, physically realistic trajectories directly from user-provided accident descriptions.
  • We demonstrate AccidentSim’s generalization across diverse road types, collision types, and driving scenarios, validating its adaptability and applicability to a wide range of real-world conditions.
2 Related Work
End-to-end models for autonomous driving perform poorly in extreme conditions, such as vehicle collision scenarios, primarily due to a lack of data for such events. Therefore, data from scenarios involving collisions plays a crucial role in enhancing the robustness of these models. To address this issue, existing approaches for acquiring and generating accident scenarios can be broadly categorized into two main strategies: real-world data collection and driving scenario generation.
Early research focused on constructing real-world accident datasets. For instance, the CADP dataset [26] provides spatio-temporal annotations of CCTV accident videos, offering valuable support for traffic accident prediction and vehicle collision detection tasks. The DADA-seg dataset [37] includes pixel-level annotations of accident scenarios, aiding models in handling extreme driving scenarios. The CCD dataset [3] further enriches accident scenario semantics by providing diverse environmental annotations and analyses of accident causes. Despite the value these real-world datasets offer in capturing collision scenarios, they face challenges such as high acquisition costs, leading to limited dataset sizes and scenario diversity.
To address the difficulty of obtaining real-world accident data, recent studies have explored various scenario generation approaches, which can be categorized into three main directions. First, language-guided traffic generation methods like CTG++ [41] and LCTGen [28] leverage large language models to control traffic simulation through natural language instructions. Second, several works focus specifically on safety-critical scenarios: CAT [39] proposes a closed-loop adversarial training framework to create challenging scenarios, while STRIVE [24] generates difficult but solvable scenarios by optimizing in the latent space of a learned traffic model. NeuroNCAP [21] advances this direction by introducing a NeRF-based simulator that simulates safety-critical scenarios through viewpoint movement for testing autonomous driving systems. Third, in the realm of video generation, methods like DriveDiffusion [19], DriveDreamer [30], and S-NeRF [34] have achieved significant progress in generating realistic driving videos, while Panacea [32], MagicDrive [12], and Drive-WM [30] have focused on improving scenario control capabilities. However, these existing methods still face limitations in simulating accurate vehicle dynamics during and after collisions that adhere to real-world physical laws, making it challenging to generate physically realistic collision scenarios.
Given the limitations of existing methods, this work introduces AccidentSim, a new framework dedicated to generating vehicle collision videos using physical cues and contextual information available in real-world accident reports, resulting in more physically realistic collision videos.
Figure 2: Architecture of AccidentSim: AccidentSim extracts physical cues, such as vehicle speeds and collision types, along with contextual information like road types, from accident reports. This extracted information is then fed into a physical simulator, such as CARLA, to generate collision trajectories, which form a dataset. The dataset is subsequently used to fine-tune a language model, AccidentLLM (based on the Llama model), enabling it to adapt to various scenarios from user-provided accident descriptions and generate novel trajectories. Building on this foundation, AccidentSim leverages user accident descriptions to plan pre-collision trajectories, generating both the pre-collision trajectory and relevant collision information, such as vehicle speeds, collision angles, and more. The collision information is then processed by AccidentLLM, which produces novel, physically realistic post-collision trajectories, resulting in complete collision trajectories. These realistic trajectories are composited onto vehicles to generate the foreground, which is then combined with rendered backgrounds to produce collision videos across various scenarios.
3 Method
AccidentSim first utilizes a large language model to extract key vehicle motion information (such as vehicle models, collision types, and intended video viewpoint movements) from the user-provided accident description. This extracted information is then processed by two core modules: the Pre-Collision Trajectory Planning module, which generates the trajectories leading to the collision, and the Post-Collision Prediction module, which generates the trajectories after the collision. Together, these modules produce a complete vehicle collision trajectory. Subsequently, in the following Collision Video Generation phase, the Foreground Processing module integrates the generated collision trajectory with user-specified vehicle models to produce foreground images. In the Background Processing module, background images are synthesized based on the user-specified viewpoint movement. Finally, the Video Composition module merges the foreground and background images to generate a vehicle collision video.
3.1 Pre-Collision Trajectory Planning
At this stage, the proposed algorithm of AccidentSim produces pre-collision trajectories leading up to the collision, along with associated information such as vehicle speed. It operates across three primary phases. First, in the Initial State Extraction phase, essential vehicle and collision parameters are extracted from the user-provided accident description to establish the initial state for each vehicle and define the collision target. Second, in the Lane and Path Selection phase, the algorithm identifies suitable lanes based on vehicle dynamics and the scene map, forming potential lane combinations to ensure feasible paths for collision trajectory planning. Third, in the Trajectory Generation and Validation phase, valid pre-collision trajectories are generated using the selected lane combinations, which are then evaluated against collision criteria to yield a final set of feasible pre-collision trajectories. For more details, refer to Section A of the Supplementary Materials.
3.1.1 Initial State Extraction
In this phase, the proposed algorithm initiates by extracting essential vehicle and collision parameters from the vehicle motion information $\mathcal{I}$, which includes the dynamics of all vehicles involved in the scenario. Using this data, the algorithm isolates variables such as the collision point $\mathcal{C}$, vehicle count $n$, distances $D$, velocities $V$, accelerations $A$, and orientation angles $\theta$. This extraction process establishes a comprehensive view of each vehicle’s current state and potential collision zones. The algorithm then determines the starting points $P_{0}$ for each vehicle by factoring in directional and dynamic properties.
3.1.2 Lane and Path Selection
This phase focuses on selecting suitable lanes within the scene map $\mathcal{M}$ to enable feasible trajectory planning. The proposed algorithm begins by parsing the scene map to identify available lanes $L$, ensuring that each candidate lane is assessed for compatibility with the vehicle’s current orientation $\theta$ and collision constraints $\mathcal{C}$. A subset of valid lanes $\mathcal{L}$ is then assembled, containing only those that meet the criteria defined by vehicle orientation and other relevant dynamics. After identifying valid lanes, the algorithm generates various lane combinations $\mathcal{G}$ by analyzing lane connectivity and spatial feasibility in relation to the parameters $n$, $D$, $V$, AA, and $\theta$.
3.1.3 Trajectory Generation and Validation
This phase involves generating and validating trajectories based on the preselected lane combinations. For each lane combination $G_{i}$ in $\mathcal{G}$, the proposed algorithm checks whether it meets specific collision criteria using the collision parameters $\mathcal{C}$. Upon identifying a valid lane combination $G_{s}$ that satisfies these conditions, it proceeds to generate the trajectory paths $\mathcal{T}$. This process incorporates the starting points $P_{0}$ and valid lane combination $G_{s}$ to form trajectories that are both dynamically feasible and collision-aware. Each trajectory is then evaluated to ensure it conforms to the criteria established for pre-collision scenarios. The final output consists of the validated set of trajectories $\mathcal{T}$ and the associated imminent collision information $\mathcal{S}$. Subsequently, the imminent collision information $\mathcal{S}$ is provided as input to our “AccidentLLM”, which predicts novel post-collision trajectories (details provided in the Post-Collision Trajectory Prediction section).
3.2 Post-Collision Trajectory Prediction
To address the challenge of collecting real-world post-collision trajectories, we leverage physical cues extracted from accident reports to train a customized version of the Llama model [11]. The resulting model, termed AccidentLLM, is capable of responding to the collision information extracted from the user-provided accident descriptions, and predicting novel post-collision trajectories that are adaptable to various road conditions.
As shown in Figure 2 (c), the proposed prediction is structured into three stages: (1) Accident Report Extraction, where collision-specific information is extracted from accident reports; (2) Collision Scenario Simulation, which generates physics-constrained post-collision trajectories; and (3) Fine-tuning AccidentLLM, where the extracted attributes and simulated scenarios are used to tailor the model for accurate trajectory prediction.
3.2.1 Accident Report Extraction
The real-world accident reports used in our work, sourced from NMVCCS[1], provide critical data for supporting collision simulations, encompassing pre-accident environmental conditions, collision dynamics, and post-accident outcomes. Prior work, such as SoVAR [14], employs a template-based approach combined with language models to extract key information from accident reports. However, due to the inherent variability and complexity of natural language descriptions in these reports, a single-template strategy lacks the flexibility to adapt to diverse accident scenarios. To overcome this limitation, we employ a Llama model [11] with a few-shot learning strategy to automatically extract essential information from accident reports, including pre-accident environmental conditions, road features, and driver behaviors; in-collision details such as vehicle positioning, speed, and collision type; and post-accident information including emergency responses. For more details, refer to Section B of the Supplementary Materials.
3.2.2 Collision Scenario Simulation
In the accident simulation phase, we developed a comprehensive scenario reconstruction pipeline to accurately recreate real-world collision scenarios and construct a high-quality collision dataset for fine-tuning AccidentLLM later. The process begins by partitioning the extracted accident report information into vehicle-specific information, such as velocity and location, and road-related details. Using RoadRunner [22], we construct high-fidelity 3D road models, which are seamlessly integrated into the CARLA simulation environment. CARLA’s advanced vehicle navigation capabilities, combined with the extracted vehicle dynamics and physical parameters such as friction and gravity, ensure realistic motion and collision behavior. Specifically, frictional forces are determined by road slope and surface properties, while gravity influences acceleration and deceleration, particularly on inclined surfaces. As the vehicle approaches the designated collision point, critical vehicle information such as speed, position, and orientation at the moment of impact are captured and the post-collision trajectory is recorded and generated. This generated post-collision trajectory data provides a robust foundation for fine-tuning the Llama model [11], significantly enhancing its ability to predict physically realistic post-collision trajectories in various scenarios. For more details, refer to Section C of the Supplementary Materials.
3.2.3 Fine-tuning AccidentLLM
After obtaining the collision dataset from the previous process, we fine-tune the Llama model [11] via LoRA [16] to specialize in predicting novel post-collision trajectories, resulting in a model we term AccidentLLM. This supervised fine-tuning optimizes the LLM for precise collision trajectory prediction in complex scenarios, responding to the user-provided accident descriptions.
To optimize AccidentLLM for precise trajectory prediction, we apply supervised fine-tuning using an $L_{1}$ loss function that measures the absolute discrepancy between predicted and ground truth trajectories across both spatial and rotational dimensions. The trajectory matching loss is defined as:
$\small L_{\text{traj}}=\frac{1}{N}\sum_{i=1}^{N}\left(\|\hat{\mathbf{p}}_{i}-\mathbf{p}_{i}^{*}\|+\|\hat{\boldsymbol{\theta}}_{i}-\boldsymbol{\theta}_{i}^{*}\|\right)$ (1)
where $\hat{\mathbf{p}}_{i}=[\hat{p}_{x,i},\hat{p}_{y,i},\hat{p}_{z,i}]^{\top}$ and $\mathbf{p}_{i}^{*}$ represent the predicted and ground-truth positions at time step $i$, respectively, while $\hat{\boldsymbol{\theta}}_{i}$ and $\boldsymbol{\theta}_{i}^{*}$ denote the predicted and true rotational vectors. $N$ is the total number of time steps in the trajectory sequence. This formulation ensures that both translational and rotational components are jointly optimized, allowing AccidentLLM to capture complex post-collision dynamics and produce physically consistent trajectory predictions in diverse crash scenarios.
3.3 Collision Video Generation
3.3.1 Foreground Processing
As shown in Figure 2 (a), the vehicle motion information extracted from the user-provided accident description undergoes Pre-Collision Trajectory Planning, which identifies the involved vehicle models and generates both the pre-collision trajectories and key collision parameters, such as vehicle speed and collision type (e.g., rear-end collision). These results are then passed to AccidentLLM, which infers physically realistic post-collision trajectories to complete the full motion paths of the vehicles involved. Based on these complete trajectories, foreground images of the physically realistic vehicle collision are rendered.
3.3.2 Background Processing
As shown in Figure 2 (b), the background processing consists of two phases: viewpoint adjustment and background rendering. In the viewpoint adjustment phase, we first parse the intended video viewpoint movements, extracted from the user-provided accident description, into positional and angular shift parameters corresponding to the target viewpoint. These shift parameters are then converted into the appropriate extrinsic transformation matrix and combined with the initial parameters to yield an updated viewpoint configuration. In the background rendering phase, if a novel viewpoint needs to be rendered, we apply the scene reconstruction method from ChatSim [31]. Otherwise, the original frame from the Waymo Open Dataset is used as the background. Specifically, ChatSim’s scene reconstruction method incorporates multi-camera alignment and luminance consistency rendering techniques, ensuring that the generated background seamlessly and naturally aligns with the vehicle’s viewpoint.
3.3.3 Video Composition
In the final video composition stage, we employ alpha channel compositing techniques to finely adjust transparency between the foreground and background images, ultimately generating a visually and physically realistic collision scenario.
Table 1: Collision rates for AccidentSim and baseline methods across diverse accident scenarios. The best value is in bold and the second-best is underlined. The results show that AccidentSim helps significantly reduce vehicle collisions by achieving lower collision rates.
Metric Methods Dual-Vehicle Scenarios Triple-Vehicle Scenarios Multi-Vehicle Scenarios Avg.
Straight Unprotected Right- Lane Straight Unprotected Right- Lane Straight Unprotected Right- Lane
Obstacle Left-Turn Turn Changing Obstacle Left-Turn Turn Changing Obstacle Left-Turn Turn Changing
Collision Rate (CR) $\downarrow$ Pre Fine-tuning 0.48 0.67 0.46 0.58 0.52 0.59 0.47 0.63 0.43 0.60 0.51 0.68 0.552
Learning-to-collide [9] 0.12 0.00 0.37 0.51 0.22 0.13 0.40 0.52 0.28 0.25 0.39 0.46 0.304
AdvSim [29] 0.23 0.05 0.41 0.53 0.31 0.10 0.38 0.45 0.28 0.17 0.35 0.58 0.320
Carla Scenario Generator [7] 0.22 0.22 0.19 0.39 0.17 0.26 0.17 0.48 0.12 0.29 0.10 0.48 0.258
Adversarial Trajectory Optimization [40] 0.14 0.00 0.23 0.30 0.21 0.04 0.28 0.21 0.23 0.02 0.31 0.18 0.179
ChatScene [38] 0.03 0.10 0.01 0.11 0.09 0.21 0.06 0.24 0.05 0.16 0.09 0.19 0.112
AccidentSim 0.05 0.08 0.01 0.09 0.07 0.11 0.13 0.12 0.02 0.07 0.10 0.03 0.073
Table 2: Quality assessment results for videos generated by AccidentSim compared to baseline methods, using VBench scores  [17], human evaluation  [6], and GPT-4o evaluation  [6]. The best value is in bold and the second-best is underlined. The results show that AccidentSim achieves state-of-the-art performance across the evaluation metrics.
Methods VBench Evaluation [17] Human Evaluation [6] GPT-4o Evaluation [6]
[…]
Kling 2.0 [18] 0.991 0.697 2.350 3.186 2.872 0.542 0.758 0.571
Gen-4 [25] 0.993 0.686 2.456 3.497 2.403 0.614 0.891 0.492
Jimeng 3.0 [4] 0.996 0.727 3.011 3.703 3.219 0.724 0.855 0.724
Cosmos [31] 0.985 0.637 2.578 3.142 2.894 0.451 0.703 0.510
AutoVFX [15] 0.990 0.701 2.983 3.286 3.450 0.521 0.768 0.682
AccidentSim 0.997 0.718 3.906 3.622 4.186 0.839 0.873 0.910
4 Experiments
To evaluate AccidentSim, we use the Waymo Open Dataset and accident reports maintained by the NHTSA, and design three experiments. First, the Collision Reduction with AccidentSim experiment demonstrates that training autonomous driving systems with collision videos generated by AccidentSim significantly reduces collision rates. Second, the Quality Evaluation for Generated Video experiment assesses both the visual fidelity and physical realism of the generated videos, demonstrating that AccidentSim achieves high quality on both fronts. Third, the Post-Collision Trajectory Evaluation experiment measures the accuracy and physical consistency of vehicle trajectories after collisions, confirming the physical plausibility and accuracy of AccidentSim’s trajectory generation. We also present qualitative results of the collision videos generated by AccidentSim, comparing them with existing methods to further highlight its superiority.
4.1 Datasets
We conducted experiments on the Waymo Open Dataset [27] and real-world accident reports maintained by the NHTSA. The Waymo Open Dataset contains data from 1,150 driving scenarios, which is suitable for covering the diverse scenarios required by AccidentSim. Each scenario in Waymo is recorded for 20 seconds and includes synchronized data from five LiDAR sensors and five high-resolution pinhole cameras. The NHTSA accident reports encompass a wide range of vehicle models, collision types, road types, and traffic conditions, making them suitable for use with AccidentSim.
Figure 3: Videos generated by AccidentSim in comparison with baseline methods. For a fair comparison, all methods are given the same user-provided accident description to depict the vehicle collision and driving scenario. Vehicles involved in the collision are boxed. The results show that AccidentSim follows the user description more accurately while achieving physically realistic collisions.
Figure 4: Vehicle collision scenarios across four typical road types. The results show that AccidentSim generates vehicle collision videos with high visual fidelity and physically realistic collision trajectories.
Table 3: The accuracy of the generated post-collision trajectories in terms of L2 (m) for AccidentSim compared to the baseline method across different accident scenarios. The results show that AccidentSim achieves overall lower average L2 errors, demonstrating more stable and precise trajectory predictions.
Accident Scenarios Methods Trajectory Prediction Error Physical Consistency Error
L2 (m) Impulse Momentum
[…]
Straight Obstacle AutoVFX 1.21 1.90 2.42 1.84 2.89 2.91 2.77 2.65
AccidentSim 1.00 1.45 2.01 1.49 2.02 1.98 2.12 2.21
Unprotected Left-Turn AutoVFX 1.02 1.43 1.82 1.42 1.93 1.87 1.76 1.84
AccidentSim 1.22 1.67 2.31 1.73 2.31 2.40 2.87 2.54
Right Turn AutoVFX 1.04 2.32 2.76 2.04 3.10 3.02 2.87 2.71
AccidentSim 0.45 0.83 1.54 0.94 0.89 0.94 0.78 0.87
Lane Changing AutoVFX 1.45 1.98 2.58 2.00 3.21 3.18 2.89 2.92
AccidentSim 0.67 1.21 1.95 1.28 1.37 1.48 1.54 1.64
4.2 Collision Reduction by AccidentSim
Collision reduction is the primary objective of using AccidentSim to train autonomous driving systems. This experiment evaluates whether the collision rate (CR) can be reduced by augmenting training with the collision videos generated by AccidentSim. A lower collision rate indicates that the generated videos more effectively improve these systems’ ability to avoid collisions.
Table 1 reports collision rates for AccidentSim across diverse accident scenarios. For baseline comparisons, on the one hand, we compare AccidentSim with several mainstream adversarial and knowledge-driven methods for generating vehicle collision scenarios in simulated environments, including Learning-to-Collide [9], AdvSim [29], the CARLA Scenario Generator [7], Adversarial Trajectory Optimization [40], and ChatScene [38]. On the other hand, we further evaluate the effect of incorporating AccidentSim’s generated collision videos on reducing collision rates. To do this, we follow the setup from ChatScene [38] and fine-tune a model trained with Soft Actor-Critic from SafeBench [35], referred to as “Pre Fine-tuning” in Table 1, without AccidentSim’s generated collision videos for training. For more details, refer to Section F of the Supplementary Materials.
The results in Table 1 show that AccidentSim helps significantly reduce vehicle collisions. Compared to the baseline models, it lowers the collision rate by up to 47.9%, outperforming all other methods in collision reduction. These findings highlight the effectiveness of incorporating collision videos generated by AccidentSim in training to enhance the robustness of autonomous driving systems.
4.3 Quality Evaluation for Generated Video
This experiment evaluates the video quality of vehicle collision scenarios generated by AccidentSim, with a focus on both visual fidelity and physical realism.
Due to the absence of ground-truth data, conventional image and video quality assessment metrics, such as PSNR, SSIM, LPIPS, FID, and FVD, are not applicable for evaluating AccidentSim-generated videos. To address this issue, we adopt the evaluation protocol of PhysGen3D [6] and conduct two complementary assessments. First, we use the mainstream benchmark suite VBench [17] to evaluate motion smoothness and imaging quality, where motion smoothness measures the smoothness of generated motion, and imaging quality assesses the presence of visual distortions that degrade perceptual quality. Second, we employ three metrics, including Physical Realism (PhysReal), Photorealism (Photoreal), and Semantic Consistency (Align), through both human evaluation and GPT-4o [23], following the evaluation pipeline of PhysGen3D [6]. Specifically, PhysReal measures the degree to which the video adheres to physical laws, capturing dynamics such as gravity, momentum, and velocity. Photoreal assesses the overall visual fidelity of the video, focusing on the presence of visual artifacts, temporal discontinuities, and the accurate depiction of lighting, shadows, and fine details. Align assesses how well the video content corresponds to the intended semantics of the user-provided accident description. As our work centers on generating collision videos from real-world accident reports, there is limited prior research that supports direct comparison. To enable a suitable evaluation, we compare AccidentSim against five state-of-the-art Image-to-Video (I2V) models that are most relevant to our task, including Kling 2.0 [18], Gen-4 [25], Jimeng 3.0 [4], Cosmos [2], and AutoVFX [15]. For a fair comparison, all models are provided with the same user’s accident description and background scenes for generating the collision videos. For more details, refer to Section E of the Supplementary Materials.
The results presented in Table 2 demonstrate that AccidentSim is capable of generating physically realistic and controllable vehicle collision videos. It achieves state-of-the-art performance in key metrics, including motion smoothness, physical realism, and semantic consistency. Concurrently, its performance in imaging quality and photorealism is highly comparable to the state-of-the-art. Furthermore, to further confirm the superiority of AccidentSim, we present qualitative results in Figure 3 to compare AccidentSim with baseline methods, and in Figure 4 to showcase generated videos across different road types.
4.4 Post-Collision Trajectory Evaluation
This experiment aims to confirm the physical plausibility and real-world consistency of post-collision vehicle trajectories generated by AccidentLLM, using quantitative comparisons against simulated ground truth.
Using vehicle collision trajectories from CARLA as ground truth, we evaluate performance based on two criteria: Trajectory Prediction Error and Physical Consistency Error. Specifically, Trajectory Prediction Error is measured by the L2 distance between predicted and ground truth trajectories. Physical Consistency Error is assessed using the Mean Absolute Error (MAE) and Root Mean Square Error (RMSE) of impulse and momentum changes before and after the collision. The physical quantities of momentum and impulse are computed with reference to the models proposed by Zhou et al. [42]. Given the limited prior work on vehicle collision trajectory prediction, we identify and adopt AutoVFX [15], which is a physics-based simulation method for generating collision videos, as the most suitable baseline due to its alignment with the physical modeling goals of our study.
As shown in Table 3, AccidentSim consistently outperforms AutoVFX in both accuracy and physical consistency. It achieves lower average L2 errors across most scenarios, particularly in Straight Obstacle, Right Turn, and Lane Changing, demonstrating more stable and precise trajectory predictions. Moreover, AccidentSim shows significantly smaller errors in impulse and momentum metrics, indicating stronger adherence to physical laws such as conservation of momentum. Additional visualizations in Figure 3 demonstrate that the simulated vehicle collisions by AccidentSim are physically constrained by gravity and exhibit smoother trajectories more consistent with real-world dynamics.
5 Conclusion
In this paper, we introduce AccidentSim, an innovative framework designed to generate vehicle collision videos with physically realistic collision trajectories. By extracting and utilizing physical cues and contextual information from real-world accident reports, AccidentSim effectively mitigates the critical scarcity of diverse collision videos essential for training robust autonomous driving models. Using a reliable physical simulation engine, AccidentSim ensures realistic vehicle motion in collision video generation. Leveraging a language model-driven approach, AccidentSim excels in predicting and planning novel, physically consistent collision trajectories. Experimental results demonstrate that AccidentSim generates vehicle collision videos with high visual fidelity and physical consistency. Future work will integrate vehicle deformation and additional physical cues to improve realism.

### v3

@@
 Future work will integrate vehicle deformation and additional physical cues to improve realism.
+Supplementary Material
+6 Table of Contents
+  • Pre-Collision Trajectory Planning
+  • Information Extraction from Accident Reports
+  • Physical Simulation of Accident Scenario
+  • Additional Qualitative Analysis of Collision Dynamics
+  • Human and GPT-4o Evaluation
+  • Collision Reduction with AccidentSim
+Appendix A Algorithm for Pre-Collision Trajectory Planning
+The proposed algorithm for pre-collision trajectory planning, mentioned in Section 3.1 of the paper, is detailed in Algorithm 1.
+Figure 5: The process of extracting physical clues and contextual information from accident reports.
+Input : Vehicle motion information $\mathcal{I}$ (includes dynamics of all vehicles); scene map $\mathcal{M}$ containing lane nodes
+Output : Valid pre-collision trajectories $\mathcal{T}$; collision information $\mathcal{S}$
+Phase 1: Initial State Extraction
+  $\{\mathcal{C},n,D,V,A,\theta\}\leftarrow\textsc{ExtractState}(\mathcal{I})$ // Extract essential vehicle and collision parameters
+  $P_{0}\leftarrow\textsc{DetermineStartingPoints}(D,\theta)$ // Determine initial positions based on dynamics and directions
+Phase 2: Lane and Path Selection
+  $L\leftarrow\textsc{ParseMap}(\mathcal{M})$ // Parse scene map to identify candidate lanes
+  $\mathcal{L}\leftarrow\emptyset$ // Initialize set for valid lanes
+  foreach lane $l_{i}\in L$ do
+     if IsValidLane($l_{i}$, $\theta$, $\mathcal{C}$) then
+        $\mathcal{L}\leftarrow\mathcal{L}\cup\{l_{i}\}$ // Add valid lane $l_{i}$ to $\mathcal{L}$
+ $\mathcal{G}\leftarrow\textsc{GatherLaneCombos}(\mathcal{L},n,D,V,A,\theta)$ // Form lane combinations for potential trajectories
+Phase 3: Trajectory Generation and Validation
+  foreach lane combination $G_{i}\in\mathcal{G}$ do
+     if MeetsCollisionCriteria($G_{i}$, $\mathcal{C}$) then
+        $G_{s}\leftarrow G_{i}$ // Select lane combination $G_{s}$ that satisfies collision criteria
+        break;
+ $\mathcal{T}\leftarrow\textsc{GenerateTrajectory}(G_{s},P_{0})$ // Generate trajectories $\mathcal{T}$ based on selected lane combo $G_{s}$, and initial points $P_{0}$
+ $\mathcal{S}\leftarrow\textsc{ImminentCollisionState}(\mathcal{T},V,A)$ // get imminent collision state $\mathcal{S}$ based on selected trajectories $\mathcal{T}$, and dynamics $\{V,A\}$
+return $\{\mathcal{T},\mathcal{S}\}$
+Algorithm 1 Pre-Collision Trajectory Planning
+Table 4: Information Extraction Accuracy: We compare AccidentSim with SoVAR[15], SoVAR_N [15], and AC3R[20] regarding their accuracy in extracting contextual information, such as road networks, traffic guidance, and dynamic object information from accident reports.
+Attributes Road Network and Traffic Guidance [15] Dynamic Objects [15]
+[…]
+SoVAR [15] 93.33% 100.00% 96.00% 76.33% 90.67% 100.00%
+SoVAR_N [15] 76.00% 91.33% 95.33% 55.00% 90.00% 100.00%
+AC3R [20] 70.67% 82.67% 78.67% 43.67% 11.67% 94.67%
+Ours 98.00% 100.00% 98.00% 93.00% 88.00% 100.00%
+Figure 6: Demonstration of physical simulations in CARLA for vehicle collisions.
+Figure 7: The collision videos generated by traditional methods such as ChatSim [34] exhibit model-interpenetration errors, where the two vehicles pass through each other.
+In contrast, AccidentSim, which incorporates physical constraints, produces physically realistic collision effects.
+Appendix B Information Extraction from Accident Reports
+The process of information extraction from the accident reports, mentioned in Section 3.2.1 of the paper, is detailed below.
+In AccidentSim, information related to the collision scene reconstruction is extracted from accident reports and analyzed using the Llama-3.1-8B model [11].
+Figure 5 shows an example of an accident report involving a collision between a Ford and a Dodge vehicle.
+In this example, AccidentSim extracts contextual information from the report, indicating that the collision occurred on a two-lane rural road with one right-curving lane in each direction and a speed limit of 89 km/h.
+The vehicles involved were a 2019 Ford F-450 Type-I ambulance and a Dodge Ram 2500 truck.
+Additionally, more detailed collision information is extracted from the report.
+For example, the report reveals that the left side of the Dodge vehicle collided with the left side of the Ford ambulance (collision type).
+At the moment of the collision, the vehicles were traveling at speeds of 49 km/h and 42 km/h, respectively.
+The collision occurred on a curve of the road (road type), with the Ford heading east and the Dodge heading west.
+The above extracted information enabled the reconstruction of road context and the replication of the collision within the CARLA simulator [10], as described in Section C of this Supplementary Material.
+We evaluated AccidentSim’s ability to extract collision information from accident reports against two mainstream methods: SoVAR [15] and AC3R [20].
+SoVAR [15] uses linguistic patterns to guide an LLM for extraction, while AC3R [20] uses NLP and hardcoded rules, requiring pre-segmented input for optimal performance.
+We also included the baseline SoVAR_N [15] in the evaluation, which directly extracts information from reports without using any linguistic patterns.
+As shown in the Table 4, AccidentSim achieves the highest accuracy for key attributes like lane number and speed limits, outperforming all the baselines.
+Furthermore, AccidentSim’s performance is also comparable to SoVAR for dynamic object information extraction.
+The superior extraction accuracy of AccidentSim is fundamental to ensure the accuracy of the subsequent simulation of vehicle collisions.
+Table 5: Comparison of Scenario Reconstruction Rate (SRR) between SoVAR and AccidentSim across different road types.
+Evaluation Metric Methods Road Types (Examples)
+Intersection T-junction Straight Road
+[…]
+AccidentSim 95.0% 76.3% 90.2%
+Appendix C Physical Simulation of Accident Scenario
+The physical simulation process for accident scenarios, mentioned in Section 3.2.2, is detailed as follows.
+As illustrated in Figure 6, we leverage key information extracted from accident reports, including physical parameters such as vehicle speed and position, as well as contextual information like road type and collision location, to conduct physical simulations of accidents within CARLA [10].
+Utilizing the extracted speed and position, we initialize vehicle states in CARLA and employ its automatic navigation capabilities to guide the vehicles toward the collision point.
+The simulation executes the collision event and captures precise data, such as the collision point and vehicle speed, via CARLA’s built-in collision sensors.
+This enables us to reproduce real-world accident scenarios with high physical fidelity, providing a reliable foundation for predicting novel collision trajectories using the fine-tuned AccidentLLM (as described in Section 3.2.3 of the paper).
+To quantitatively evaluate the effectiveness of AccidentSim in reconstructing accident scenarios, we adopted the methodology from SoVAR [15], using the Scenario Reconstruction Rate (SRR) as the primary evaluation metric.
+The SRR measures the overall proportion of trajectories generated from accident reports that successfully reproduce the reported collisions.
+Consequently, a higher SRR indicates superior accuracy in reconstructing real-world accident scenarios.
+As shown in Table 5, the results demonstrate that AccidentSim achieves higher reproduction accuracy than SoVAR across a range of randomly selected road configurations.
+Notably, the improvements are particularly significant for road types of T-intersections and straight roads.
+We attribute these advancements to AccidentSim’s enhanced trajectory planning capabilities resulting from the incorporation of physical constraints, which leads to greater accuracy and controllability in complex scenarios.
+Appendix D Additional Qualitative Analysis of Collision Dynamics
+Existing generative methods often struggle to produce physically realistic collision behaviors due to the lack of explicit physical constraints.
+To demonstrate and evaluate this, we compare AccidentSim with ChatSim [34], the first system that enables editable 3D driving scenes via natural language commands and external digital assets.
+As illustrated in Figure 7, ChatSim tends to exhibit severe errors of model interpenetration (i.e., vehicles passing through each other) during the collision.
+In contrast, AccidentSim leverages physical simulation to enforce rigid body dynamics and strict collision boundaries.
+The frame-by-frame analysis confirms that AccidentSim effectively eliminates such visual errors, yielding physically plausible outcomes that are consistent with real-world dynamics.
+Appendix E Human and GPT-4o Evaluation
+Figure 8: Prompt used for GPT-4o evaluation.
+Figure 9: Results of human evaluation.
+E.1 Human Evaluation
+The process of the human evaluation, mentioned in Section 4.3 of the paper, is detailed below.
+We recruited 30 participants aged 23 to 35 years, with a standard deviation of 3.70 years.
+The cohort comprised 20 males and 10 females, of whom 18 had at least two years of driving experience.
+Participants represented a range of professional backgrounds, including students, software professionals, financial sector employees, and transportation industry workers.
+We generated a total of 72 videos representing diverse vehicle collision scenarios, utilizing 6 methods as described in the paper, each covering 6 distinct collision types and 6 different scenarios.
+Each participant was requested to view and evaluate the generated collision videos.
+Participants first viewed each video and then completed a questionnaire, rating Physical Realism, Photorealism, and Semantic Consistency.
+To ensure the fairness and consistency of the evaluations, we provided detailed scoring criteria in the questionnaire, enabling participants to make informed and targeted assessments.
+The detailed evaluation results are presented in Figure 9.
+The experimental results show that AccidentSim achieves the best performance in terms of physical realism and semantic consistency, while being highly comparable to the state-of-the-art in photorealism.
+E.2 GPT-4o Evaluation
+The process of the GPT-4o evaluation, mentioned in Section 4.3 of the paper, is detailed below.
+We followed PhysGen3D [6] and employed GPT-4o as a multimodal evaluator to quantitatively assess the quality of the generated vehicle collision videos.
+This AI-based evaluation serves as an automated counterpart to human judgment, offering an additional layer of validation.
+Similar to the human evaluation, the evaluation centers on three core dimensions: Physical Realism, Photorealism, and Semantic Consistency.
+Each video is processed using a structured prompt specifically designed for GPT-4o.
+The prompt instructs the model to assign numerical scores from 1 to 5 to each of the three dimensions.
+The detailed definitions of these scores are aligned with the criteria used in the human evaluation to ensure consistency across assessment modalities.
+Figure 8 provides the prompt words we used when testing with GPT-4o.
+Appendix F Collision Reduction with AccidentSim
+For the experiment on collision reduction by AccidentSim, mentioned in Section 4.2 of the paper, the experimental configuration and training details are provided below.
+We conduct our experiments on the SafeBench platform [38] with an Nvidia A40 GPU.
+For scenario generation, we transfer vehicle driving situations from complex accidents reported by NHTSA to the driving scenarios in the Waymo Open Dataset and use the Scenic language [12] to script and simulate accident scenarios in Waymo environments within CARLA.
+In total, 4000 training scenarios and 1000 testing scenarios are generated.
+For the classification of accident scenarios, we adopt the accident scenario categorization proposed in ChatScene [41] and extend it to include three-vehicle and multi-vehicle collision scenarios.
+For training, we employ the SAC algorithm [16] to train the autonomous vehicle’s control policy.
+The SAC agent is trained for 500 epochs with a learning rate of 0.0001, and we evaluate its performance every 50 epochs to report the optimal results.

### v4

@@
 These trajectories are then used to fine-tune a language model, termed AccidentLLM, enabling it to predict novel, physically consistent trajectories directly from the collision information extracted from the user-provided accident descriptions, extending beyond the collision scenarios contained in the original accident reports.
-1 Introduction
+I Introduction
 Research and development in autonomous driving critically depend on access to diverse and comprehensive training data to ensure robust performance across various driving scenarios.
@@
 To address these challenges, we propose AccidentSim, a framework that introduces a novel approach to generating physically realistic vehicle collision videos based on real-world accident reports.
-For instance, accident reports from the National Motor Vehicle Crash Causation Survey (NMVCCS) [1] in the United States, document actual vehicle collisions across diverse vehicle models, collision types, road types, and traffic conditions.
-AccidentSim leverages the rich physical cues and contextual information embedded in such reports to synthesize vehicle-collision videos with physically realistic post-collision trajectories.
+For instance, accident reports from the National Motor Vehicle Crash Causation Survey (NMVCCS) [34] in the United States document actual vehicle collisions across diverse vehicle models, collision types, road types, and traffic conditions.
+AccidentSim leverages the rich physical cues and contextual information embedded in such reports to synthesize vehicle collision videos with physically realistic post-collision trajectories.
 As illustrated in Figure 1, AccidentSim first extracts key physical cues, such as vehicle speed, collision angle, and mass, together with contextual information, including road type and collision type, from the accident reports.
@@
   • We demonstrate AccidentSim’s generalization across diverse road types, collision types, and driving scenarios, validating its adaptability and applicability to a wide range of real-world conditions.
-2 Related Work
-End-to-end models for autonomous driving perform poorly in extreme conditions, such as vehicle collision scenarios, primarily due to a lack of data for such events.
-Therefore, data from scenarios involving collisions plays a crucial role in enhancing the robustness of these models.
-To address this issue, existing approaches for acquiring and generating accident scenarios can be broadly categorized into two main strategies: real-world data collection and driving scenario generation.
-Early research focused on constructing real-world accident datasets.
-For instance, the CADP dataset [29] provides spatio-temporal annotations of CCTV accident videos, offering valuable support for traffic accident prediction and vehicle collision detection tasks.
-The DADA-seg dataset [40] includes pixel-level annotations of accident scenarios, aiding models in handling extreme driving scenarios.
-The CCD dataset [3] further enriches accident scenario semantics by providing diverse environmental annotations and analyses of accident causes.
-Despite the value these real-world datasets offer in capturing collision scenarios, they face challenges such as high acquisition costs, leading to limited dataset sizes and scenario diversity.
-To address the difficulty of obtaining real-world accident data, recent studies have explored various scenario generation approaches, which can be categorized into three main directions.
-First, language-guided traffic generation methods like CTG++ [44] and LCTGen [31] leverage large language models to control traffic simulation through natural language instructions.
-Second, several works focus specifically on safety-critical scenarios: CAT [42] proposes a closed-loop adversarial training framework to create challenging scenarios, while STRIVE [27] generates difficult but solvable scenarios by optimizing in the latent space of a learned traffic model.
-NeuroNCAP [24] advances this direction by introducing a NeRF-based simulator that simulates safety-critical scenarios through viewpoint movement for testing autonomous driving systems.
-Third, in the realm of video generation, methods like DriveDiffusion [22], DriveDreamer [33], and S-NeRF [37] have achieved significant progress in generating realistic driving videos, while Panacea [35], MagicDrive [13], and Drive-WM [33] have focused on improving scenario control capabilities.
-However, these existing methods still face limitations in simulating accurate vehicle dynamics during and after collisions that adhere to real-world physical laws, making it challenging to generate physically realistic collision scenarios.
-Given the limitations of existing methods, this work introduces AccidentSim, a new framework dedicated to generating vehicle collision videos using physical cues and contextual information available in real-world accident reports, resulting in more physically realistic collision videos.
+II Related Work
+II-A Safety-Critical Scenarios and Driving Datasets
+The foundation of autonomous driving research relies on large-scale datasets, such as KITTI [17], Cityscapes [8], nuScenes [4], Waymo Open Dataset [44], and BDD100K [61].
+While providing massive normal driving trajectories, they inherently lack safety-critical vehicle collisions.
+To address this, early efforts constructed specialized accident datasets, such as CADP [42], ISSAFE [63], and CCD [2], which are enriched with spatio-temporal and environmental annotations.
+Recently, advanced paradigms like EQ-TAA [12] employ video diffusion to synthesize pseudo-accident pairs to actively mitigate data bias.
+However, these datasets and frameworks are bottlenecked by fixed viewpoints and prohibitive acquisition costs.
+More importantly, while excelling at observational anticipation, they fundamentally lack the capability to dynamically synthesize novel crash scenarios.
+To bypass real-world data collection, adversarial scenario generation and trajectory optimization have been widely explored.
+Methods like AdvSim [47], Learning-to-Collide [9], and CAT [65] employ adversarial training for challenging pre-collision scenarios, while STRIVE [38] and others [66] optimize latent traffic models evaluated in platforms like SafeBench [55].
+However, these approaches focus solely on pre-collision obstacle avoidance.
+They fail to simulate actual physical impacts and post-collision dynamics, leaving a critical gap in fully replicating the accident lifecycle.
+II-B Trajectory Prediction and Language-Guided Generation
+Accurate vehicle motion forecasting is core to driving simulation.
+Traditional multi-agent trajectory prediction models [43, 25, 35, 13] rely heavily on historical states, latent sequence modeling [18], and high-definition maps.
+To explicitly capture temporal dependencies across multiple trajectories, recent efforts integrate spatio-temporal sequence fusion to comprehensively model social interactions among agents [48].
+Recently, language-guided traffic generation driven by large language models (LLMs) [11] has emerged.
+Frameworks like LCTGen [45], CTG++ [68], TrafficGen [14], and Traj-LLM [58] control traffic simulation via natural language, while MotionLM [41] and MotionCLIP [46] frame forecasting directly as a language modeling task.
+To bridge high-level reasoning and low-level control, recent studies distill multi-modal LLM knowledge into efficient vision-based planners [20], enhancing generalizability in critical scenarios without prohibitive inference costs.
+Alternatively, to manage online querying overhead, AdaDrive [67] proposes an adaptive collaborative system.
+By dynamically invoking the LLM exclusively during complex scenarios, it balances high-level reasoning with fast conventional planners.
+II-C Physical Simulation and Collision Modeling
+Early analytical models [37, 69] utilized momentum, restitution, and friction to predict post-impact motion.
+Modern research integrates these principles into high-fidelity 3D simulators like CARLA [10], LGSVL [39], and UniSim [59] to enforce strict physical constraints.
+While these engines provide mathematically rigorous forward kinematics, integrating them into data-driven generative pipelines remains challenging.
+Recently, SoVAR [19] and AC3R [24] parsed real-world police reports from NMVCCS [34] to reconstruct accidents in simulators.
+Concurrently, generative architectures—such as combining Variational Autoencoders with Wasserstein GANs (VAE-WGAN)—have been integrated into end-to-end Deep Reinforcement Learning (DRL) frameworks to efficiently encode video features and explicitly minimize vehicle collision rates in simulators like CARLA [26].
+However, pure physical simulations are computationally expensive, require meticulous parameter tuning, and cannot dynamically generate novel scenarios from under-specified prompts.
+AccidentSim resolves this by using simulators solely to construct a high-quality dataset.
+Distilling this data into AccidentLLM enables real-time, physics-aligned trajectory generation without the overhead of online physical rendering.
+II-D Controllable Video Generation and World Models
+Diffusion models have profoundly transformed driving simulation.
+Foundation models (e.g., Kling 2.0 [27], Gen-4 [40], Jimeng 3.0 [3], CogVideoX [60]) establish new photorealism benchmarks.
+Tailored for autonomous driving, generative world models [49, 28, 16, 51, 53, 33, 1] synthesize videos conditioned on layouts, text, or 3D geometry.
+Concurrently, NeRF and 3DGS advancements [54, 64, 31, 50, 70, 57] enable editable large-scale dynamic scene simulation.
+To further enhance 3D perception and dataset diversity, hardness-aware scene synthesis [62] adaptively composes 3D objects and backgrounds.
+Moving beyond normal conditions, frameworks like Challenger [56] utilize physics-aware trajectory refinement to synthesize aggressive adversarial scenarios (e.g., cut-ins).
+To mitigate computational overhead in high-resolution multi-view synthesis, DriveScape [52] efficiently fuses multi-view features with 3D road structures.
+Moreover, self-consistent generative networks [30] rigidly preserve spatial consistency across diverse viewpoints with large camera baselines, bypassing explicit geometric rectification.
+Finally, MagicDrive-V2 [15] achieves high-resolution, long video generation with precise geometric control, while recent disentangled fine-tuning strategies [6] prevent attribute confusion by accurately binding specific motions and visual concepts to individual interacting agents in complex multi-vehicle scenarios.
 Figure 2: Architecture of AccidentSim: AccidentSim extracts physical cues, such as vehicle speeds and collision types, along with contextual information like road types, from accident reports.
@@
 These realistic trajectories are composited onto vehicles to generate the foreground, which is then combined with rendered backgrounds to produce collision videos across various scenarios.
-3 Method
+III METHODOLOGY
 AccidentSim first utilizes a large language model to extract key vehicle motion information (such as vehicle models, collision types, and intended video viewpoint movements) from the user-provided accident description.
@@
 Finally, the Video Composition module merges the foreground and background images to generate a vehicle collision video.
-3.1 Pre-Collision Trajectory Planning
+III-A Pre-Collision Trajectory Planning
 At this stage, the proposed algorithm of AccidentSim produces pre-collision trajectories leading up to the collision, along with associated information such as vehicle speed.
@@
 Third, in the Trajectory Generation and Validation phase, valid pre-collision trajectories are generated using the selected lane combinations, which are then evaluated against collision criteria to yield a final set of feasible pre-collision trajectories.
-For more details, refer to Section A of the Supplementary Materials.
-3.1.1 Initial State Extraction
+III-A1 Initial State Extraction
 In this phase, the proposed algorithm initiates by extracting essential vehicle and collision parameters from the vehicle motion information $\mathcal{I}$, which includes the dynamics of all vehicles involved in the scenario.
@@
 The algorithm then determines the starting points $P_{0}$ for each vehicle by factoring in directional and dynamic properties.
-3.1.2 Lane and Path Selection
+III-A2 Lane and Path Selection
 This phase focuses on selecting suitable lanes within the scene map $\mathcal{M}$ to enable feasible trajectory planning.
@@
 After identifying valid lanes, the algorithm generates various lane combinations $\mathcal{G}$ by analyzing lane connectivity and spatial feasibility in relation to the parameters $n$, $D$, $V$, AA, and $\theta$.
-3.1.3 Trajectory Generation and Validation
+III-A3 Trajectory Generation and Validation
 This phase involves generating and validating trajectories based on the preselected lane combinations.
@@
 Subsequently, the imminent collision information $\mathcal{S}$ is provided as input to our “AccidentLLM”, which predicts novel post-collision trajectories (details provided in the Post-Collision Trajectory Prediction section).
-3.2 Post-Collision Trajectory Prediction
+III-B Post-Collision Trajectory Prediction
 To address the challenge of collecting real-world post-collision trajectories, we leverage physical cues extracted from accident reports to train a customized version of the Llama model [11].
@@
 As shown in Figure 2 (c), the proposed prediction is structured into three stages: (1) Accident Report Extraction, where collision-specific information is extracted from accident reports; (2) Collision Scenario Simulation, which generates physics-constrained post-collision trajectories; and (3) Fine-tuning AccidentLLM, where the extracted attributes and simulated scenarios are used to tailor the model for accurate trajectory prediction.
-3.2.1 Accident Report Extraction
+III-B1 Accident Report Extraction
 The real-world accident reports used in our work, sourced from NMVCCS [34], provide critical data for supporting collision simulations, encompassing pre-accident environmental conditions, collision dynamics, and post-accident outcomes.
@@
 To overcome this limitation, we employ a Llama model [11] with a few-shot learning strategy to automatically extract essential information from accident reports, including pre-accident environmental conditions, road features, and driver behaviors; in-collision details such as vehicle positioning, speed, and collision type; and post-accident information including emergency responses.
-For more details, refer to Section B of the Supplementary Materials.
-3.2.2 Collision Scenario Simulation
-In the accident simulation phase, we developed a comprehensive scenario reconstruction pipeline to accurately recreate real-world collision scenarios and construct a high-quality collision dataset for fine-tuning AccidentLLM later.
+III-B2 Collision Scenario Simulation
+In the accident simulation phase, we developed a comprehensive scenario reconstruction pipeline to accurately recreate real-world collision scenarios and construct a high-quality collision dataset for the subsequent fine-tuning of AccidentLLM.
 The process begins by partitioning the extracted accident report information into vehicle-specific information, such as velocity and location, and road-related details.
@@
 This generated post-collision trajectory data provides a robust foundation for fine-tuning the Llama model [11], significantly enhancing its ability to predict physically realistic post-collision trajectories in various scenarios.
-For more details, refer to Section C of the Supplementary Materials.
-3.2.3 Fine-tuning AccidentLLM
+III-B3 Fine-tuning AccidentLLM
 After obtaining the collision dataset from the previous process, we fine-tune the Llama model [11] via LoRA [22] to specialize in predicting novel post-collision trajectories, resulting in a model we term AccidentLLM.
@@
 The trajectory matching loss is defined as:
-$\small L_{\text{traj}}=\frac{1}{N}\sum_{i=1}^{N}\left(\|\hat{\mathbf{p}}_{i}-\mathbf{p}_{i}^{*}\|+\|\hat{\boldsymbol{\theta}}_{i}-\boldsymbol{\theta}_{i}^{*}\|\right)$ (1)
+$L_{\text{traj}}=\frac{1}{N}\sum_{i=1}^{N}\left(\|\hat{\mathbf{p}}_{i}-\mathbf{p}_{i}^{*}\|+\|\hat{\boldsymbol{\theta}}_{i}-\boldsymbol{\theta}_{i}^{*}\|\right)$ (1)
 where $\hat{\mathbf{p}}_{i}=[\hat{p}_{x,i},\hat{p}_{y,i},\hat{p}_{z,i}]^{\top}$ and $\mathbf{p}_{i}^{*}$ represent the predicted and ground-truth positions at time step $i$, respectively, while $\hat{\boldsymbol{\theta}}_{i}$ and $\boldsymbol{\theta}_{i}^{*}$ denote the predicted and true rotational vectors.
@@
 This formulation ensures that both translational and rotational components are jointly optimized, allowing AccidentLLM to capture complex post-collision dynamics and produce physically consistent trajectory predictions in diverse crash scenarios.
-3.3 Collision Video Generation
-3.3.1 Foreground Processing
+III-C Collision Video Generation
+III-C1 Foreground Processing
 As shown in Figure 2 (a), the vehicle motion information extracted from the user-provided accident description undergoes Pre-Collision Trajectory Planning, which identifies the involved vehicle models and generates both the pre-collision trajectories and key collision parameters, such as vehicle speed and collision type (e.g., rear-end collision).
@@
 Based on these complete trajectories, foreground images of the physically realistic vehicle collision are rendered.
-3.3.2 Background Processing
+III-C2 Background Processing
 As shown in Figure 2 (b), the background processing consists of two phases: viewpoint adjustment and background rendering.
@@
 Specifically, ChatSim’s scene reconstruction method incorporates multi-camera alignment and luminance consistency rendering techniques, ensuring that the generated background seamlessly and naturally aligns with the vehicle’s viewpoint.
-3.3.3 Video Composition
+III-C3 Video Composition
 In the final video composition stage, we employ alpha channel compositing techniques to finely adjust transparency between the foreground and background images, ultimately generating a visually and physically realistic collision scenario.
-Table 1: Collision rates for AccidentSim and baseline methods across diverse accident scenarios.
-The best value is in bold and the second-best is underlined.
-The results show that AccidentSim helps significantly reduce vehicle collisions by achieving lower collision rates.
-Metric Methods Dual-Vehicle Scenarios Triple-Vehicle Scenarios Multi-Vehicle Scenarios Avg.
-Straight Unprotected Right- Lane Straight Unprotected Right- Lane Straight Unprotected Right- Lane
-Obstacle Left-Turn Turn Changing Obstacle Left-Turn Turn Changing Obstacle Left-Turn Turn Changing
-Collision Rate (CR) $\downarrow$ Pre Fine-tuning 0.48 0.67 0.46 0.58 0.52 0.59 0.47 0.63 0.43 0.60 0.51 0.68 0.552
-Learning-to-collide [9] 0.12 0.00 0.37 0.51 0.22 0.13 0.40 0.52 0.28 0.25 0.39 0.46 0.304
-AdvSim [32] 0.23 0.05 0.41 0.53 0.31 0.10 0.38 0.45 0.28 0.17 0.35 0.58 0.320
-Carla Scenario Generator [7] 0.22 0.22 0.19 0.39 0.17 0.26 0.17 0.48 0.12 0.29 0.10 0.48 0.258
-Adversarial Trajectory Optimization [43] 0.14 0.00 0.23 0.30 0.21 0.04 0.28 0.21 0.23 0.02 0.31 0.18 0.179
-ChatScene [41] 0.03 0.10 0.01 0.11 0.09 0.21 0.06 0.24 0.05 0.16 0.09 0.19 0.112
-AccidentSim 0.05 0.08 0.01 0.09 0.07 0.11 0.13 0.12 0.02 0.07 0.10 0.03 0.073
-Table 2: Quality assessment results for videos generated by AccidentSim compared to baseline methods, using VBench scores  [19], human evaluation  [6], and GPT-4o evaluation  [6].
+IV Experiments
+To evaluate AccidentSim, we utilize the Waymo Open Dataset and NHTSA accident reports (Section IV-A).
+Our experiments focus on three core aspects: (1) Trajectory Generation Quality and Physical Consistency (Section IV-B) confirms the kinematic accuracy and semantic alignment of the generated post-collision trajectories.
+(2) Quality Evaluation for Generated Video (Section IV-C) demonstrates AccidentSim’s superior visual fidelity and physical realism over existing baselines.
+(3) Collision Reduction by AccidentSim (Section IV-D) validates the practical utility of our synthesized videos in significantly reducing collision rates for autonomous driving systems.
+IV-A Datasets
+We conducted experiments on the Waymo Open Dataset [44] and real-world accident reports maintained by the NHTSA.
+The Waymo Open Dataset contains data from 1,150 driving scenarios, which is suitable for covering the diverse scenarios required by AccidentSim.
+Each scenario in Waymo is recorded for 20 seconds and includes synchronized data from five LiDAR sensors and five high-resolution pinhole cameras.
+The NHTSA accident reports encompass a wide range of vehicle models, collision types, road types, and traffic conditions, making them suitable for use with AccidentSim.
+Figure 3: Videos generated by AccidentSim in comparison with baseline methods.
+For a fair comparison, all methods are given the same user-provided accident description to depict the vehicle collision and driving scenario.
+Vehicles involved in the collision are boxed.
+The results show that AccidentSim follows the user description more accurately while achieving physically realistic collisions.
+TABLE I: The accuracy of the generated post-collision trajectories in terms of L2 (m) for AccidentSim compared to the baseline method across different accident scenarios.
+The results show that AccidentSim achieves overall lower average L2 errors, demonstrating more stable and precise trajectory predictions.
+Accident Scenarios Methods Trajectory Prediction Error Physical Consistency Error
+L2 (m) Impulse Momentum
+[…]
+AutoVFX 1.21 1.90 2.42 1.84 2.89 2.91 2.77 2.65
+[…]
+AutoVFX 1.02 1.43 1.82 1.42 1.93 1.87 1.76 1.84
+[…]
+AutoVFX 1.04 2.32 2.76 2.04 3.10 3.02 2.87 2.71
+[…]
+AutoVFX 1.45 1.98 2.58 2.00 3.21 3.18 2.89 2.92
+[…]
+TABLE II: Human study results on the language-conditioned simulation.
+Methods Crash Report Attribute Description
+Ours Preferred (%) $\uparrow$ Score (1–5) $\uparrow$ Ours Preferred (%) $\uparrow$ Score (1–5) $\uparrow$
+TrafficGen [14] 95.68 1.49 94.51 2.63
+MotionCLIP [46] 96.49 1.71 95.71 2.20
+LCTGen [45] 89.53 3.93 87.29 4.10
+[…]
+TABLE III: Quality assessment results for videos generated by AccidentSim compared to baseline methods, using VBench scores  [23], human evaluation  [5], and GPT-4o evaluation  [5].
 The best value is in bold and the second-best is underlined.
@@
 AutoVFX [21] 0.990 0.701 2.983 3.286 3.450 0.521 0.768 0.682
-AccidentSim 0.997 0.718 3.906 3.622 4.186 0.839 0.873 0.910
-4 Experiments
-To evaluate AccidentSim, we use the Waymo Open Dataset and accident reports maintained by the NHTSA, and design three experiments.
-First, the Collision Reduction with AccidentSim experiment demonstrates that training autonomous driving systems with collision videos generated by AccidentSim significantly reduces collision rates.
-Second, the Quality Evaluation for Generated Video experiment assesses both the visual fidelity and physical realism of the generated videos, demonstrating that AccidentSim achieves high quality on both fronts.
-Third, the Post-Collision Trajectory Evaluation experiment measures the accuracy and physical consistency of vehicle trajectories after collisions, confirming the physical plausibility and accuracy of AccidentSim’s trajectory generation.
-We also present qualitative results of the collision videos generated by AccidentSim, comparing them with existing methods to further highlight its superiority.
-4.1 Datasets
-We conducted experiments on the Waymo Open Dataset [30] and real-world accident reports maintained by the NHTSA.
-The Waymo Open Dataset contains data from 1,150 driving scenarios, which is suitable for covering the diverse scenarios required by AccidentSim.
-Each scenario in Waymo is recorded for 20 seconds and includes synchronized data from five LiDAR sensors and five high-resolution pinhole cameras.
-The NHTSA accident reports encompass a wide range of vehicle models, collision types, road types, and traffic conditions, making them suitable for use with AccidentSim.
-Figure 3: Videos generated by AccidentSim in comparison with baseline methods.
-For a fair comparison, all methods are given the same user-provided accident description to depict the vehicle collision and driving scenario.
-Vehicles involved in the collision are boxed.
-The results show that AccidentSim follows the user description more accurately while achieving physically realistic collisions.
+[…]
+TABLE IV: Collision rates for AccidentSim and baseline methods across diverse accident scenarios.
+The best value is in bold and the second-best is underlined.
+The results show that AccidentSim helps significantly reduce vehicle collisions by achieving lower collision rates.
+Methods Dual-Vehicle Scenarios Triple-Vehicle Scenarios Multi-Vehicle Scenarios Avg.
+Straight Obstacle Unprotected Left-Turn Right- Turn Lane Changing Straight Obstacle Unprotected Left-Turn Right- Turn Lane Changing Straight Obstacle Unprotected Left-Turn Right- Turn Lane Changing
+Pre Fine-tuning 0.48 0.67 0.46 0.58 0.52 0.59 0.47 0.63 0.43 0.60 0.51 0.68 0.552
+Learning-to-collide [9] 0.12 0.00 0.37 0.51 0.22 0.13 0.40 0.52 0.28 0.25 0.39 0.46 0.304
+AdvSim [47] 0.23 0.05 0.41 0.53 0.31 0.10 0.38 0.45 0.28 0.17 0.35 0.58 0.320
+Carla Scenario Generator [7] 0.22 0.22 0.19 0.39 0.17 0.26 0.17 0.48 0.12 0.29 0.10 0.48 0.258
+Adversarial Trajectory Optimization [66] 0.14 0.00 0.23 0.30 0.21 0.04 0.28 0.21 0.23 0.02 0.31 0.18 0.179
+ChatScene [64] 0.03 0.10 0.01 0.11 0.09 0.21 0.06 0.24 0.05 0.16 0.09 0.19 0.112
+[…]
 Figure 4: Vehicle collision scenarios across four typical road types.
 The results show that AccidentSim generates vehicle collision videos with high visual fidelity and physically realistic collision trajectories.
-Table 3: The accuracy of the generated post-collision trajectories in terms of L2 (m) for AccidentSim compared to the baseline method across different accident scenarios.
-The results show that AccidentSim achieves overall lower average L2 errors, demonstrating more stable and precise trajectory predictions.
-Accident Scenarios Methods Trajectory Prediction Error Physical Consistency Error
-L2 (m) Impulse Momentum
-[…]
-Straight Obstacle AutoVFX 1.21 1.90 2.42 1.84 2.89 2.91 2.77 2.65
-AccidentSim 1.00 1.45 2.01 1.49 2.02 1.98 2.12 2.21
-Unprotected Left-Turn AutoVFX 1.02 1.43 1.82 1.42 1.93 1.87 1.76 1.84
-AccidentSim 1.22 1.67 2.31 1.73 2.31 2.40 2.87 2.54
-Right Turn AutoVFX 1.04 2.32 2.76 2.04 3.10 3.02 2.87 2.71
-AccidentSim 0.45 0.83 1.54 0.94 0.89 0.94 0.78 0.87
-Lane Changing AutoVFX 1.45 1.98 2.58 2.00 3.21 3.18 2.89 2.92
-AccidentSim 0.67 1.21 1.95 1.28 1.37 1.48 1.54 1.64
-4.2 Collision Reduction by AccidentSim
+IV-B Trajectory Generation Quality and Physical Consistency
+1) Physical Consistency and Accuracy: We evaluate the physical plausibility of AccidentLLM’s post-collision trajectories against CARLA-simulated ground truth.
+Trajectory Prediction Error is measured via L2 distance, and Physical Consistency Error is assessed using the MAE and RMSE of impulse and momentum changes [69].
+We select AutoVFX [21], a physics-based simulation method, as our primary baseline.
+As shown in Table I, AccidentSim consistently outperforms AutoVFX.
+It achieves lower average L2 errors and significantly smaller impulse and momentum errors, demonstrating stricter adherence to physical laws like momentum conservation.
+Figure 3 further visualizes these physically constrained trajectories, which exhibit smoother and more realistic dynamics.
+2) Text-to-Trajectory Semantic Consistency: Following the LCTGen [45] protocol, we conducted human A/B tests to measure Preference (Ours Preferred %) and Alignment (1–5 scores) on Crash Report and Attribute Description datasets.
+Table II demonstrates that AccidentSim significantly outperforms baselines.
+By fine-tuning on simulator-calibrated data, AccidentLLM effectively maps complex textual narratives to kinematically consistent trajectories.
+IV-C Quality Evaluation for Generated Video
+This experiment evaluates the visual fidelity and physical realism of generated vehicle collisions.
+Since conventional metrics (e.g., PSNR, FID, FVD) require ground-truth data, we adopt the PhysGen3D [5] protocol for reference-free evaluation.
+First, we use VBench [23] to assess motion smoothness and imaging quality.
+Second, we employ human and GPT-4o [36] evaluations to measure three key metrics: Physical Realism (PhysReal), Photorealism (PhotoReal), and Semantic Consistency (Align).
+Informed consent was obtained from all participants involved in the human evaluation.
+Specifically, PhysReal measures adherence to physical laws, capturing dynamics like gravity and momentum; PhotoReal assesses overall visual fidelity, focusing on the absence of artifacts and the accurate depiction of lighting and shadows; and Align evaluates how accurately the generated video reflects the semantics of the user-provided accident description.
+We compare AccidentSim against five state-of-the-art video generation models: Kling 2.0 [27], Gen-4 [40], Jimeng 3.0 [3], Cosmos [1], and AutoVFX [21].
+For a fair comparison, all models receive identical accident descriptions and background scenes.
+As shown in Table III, AccidentSim achieves state-of-the-art performance in motion smoothness, physical realism, and semantic consistency, while maintaining highly competitive imaging quality.
+Qualitative comparisons in Figure 3 and diverse scenario results in Figure 4 further demonstrate AccidentSim’s superiority in generating physically realistic and controllable vehicle collisions.
+IV-D Collision Reduction by AccidentSim
 Collision reduction is the primary objective of using AccidentSim to train autonomous driving systems.
@@
 A lower collision rate indicates that the generated videos more effectively improve these systems’ ability to avoid collisions.
-Table 1 reports collision rates for AccidentSim across diverse accident scenarios.
+Table IV reports collision rates for AccidentSim across diverse accident scenarios.
 For baseline comparisons, on the one hand, we compare AccidentSim with several mainstream adversarial and knowledge-driven methods for generating vehicle collision scenarios in simulated environments, including Learning-to-Collide [9], AdvSim [47], the CARLA Scenario Generator [7], Adversarial Trajectory Optimization [66], and ChatScene [64].
 On the other hand, we further evaluate the effect of incorporating AccidentSim’s generated collision videos on reducing collision rates.
-To do this, we follow the setup from ChatScene [41] and fine-tune a model trained with Soft Actor-Critic from SafeBench [38], referred to as “Pre Fine-tuning” in Table 1, without AccidentSim’s generated collision videos for training.
-For more details, refer to Section F of the Supplementary Materials.
-The results in Table 1 show that AccidentSim helps significantly reduce vehicle collisions.
-Compared to the baseline models, it lowers the collision rate by up to 47.9%, outperforming all other methods in collision reduction.
-These findings highlight the effectiveness of incorporating collision videos generated by AccidentSim in training to enhance the robustness of autonomous driving systems.
-4.3 Quality Evaluation for Generated Video
-This experiment evaluates the video quality of vehicle collision scenarios generated by AccidentSim, with a focus on both visual fidelity and physical realism.
-Due to the absence of ground-truth data, conventional image and video quality assessment metrics, such as PSNR, SSIM, LPIPS, FID, and FVD, are not applicable for evaluating AccidentSim-generated videos.
-To address this issue, we adopt the evaluation protocol of PhysGen3D [6] and conduct two complementary assessments.
-First, we use the mainstream benchmark suite VBench [19] to evaluate motion smoothness and imaging quality, where motion smoothness measures the smoothness of generated motion, and imaging quality assesses the presence of visual distortions that degrade perceptual quality.
-Second, we employ three metrics, including Physical Realism (PhysReal), Photorealism (Photoreal), and Semantic Consistency (Align), through both human evaluation and GPT-4o [26], following the evaluation pipeline of PhysGen3D [6].
-Specifically, PhysReal measures the degree to which the video adheres to physical laws, capturing dynamics such as gravity, momentum, and velocity.
-Photoreal assesses the overall visual fidelity of the video, focusing on the presence of visual artifacts, temporal discontinuities, and the accurate depiction of lighting, shadows, and fine details.
-Align assesses how well the video content corresponds to the intended semantics of the user-provided accident description.
-As our work centers on generating collision videos from real-world accident reports, there is limited prior research that supports direct comparison.
-To enable a suitable evaluation, we compare AccidentSim against five state-of-the-art Image-to-Video (I2V) models that are most relevant to our task, including Kling 2.0 [21], Gen-4 [28], Jimeng 3.0 [4], Cosmos [2], and AutoVFX [17].
-For a fair comparison, all models are provided with the same user’s accident description and background scenes for generating the collision videos.
-For more details, refer to Section E of the Supplementary Materials.
-The results presented in Table 2 demonstrate that AccidentSim is capable of generating physically realistic and controllable vehicle collision videos.
-It achieves state-of-the-art performance in key metrics, including motion smoothness, physical realism, and semantic consistency.
-Concurrently, its performance in imaging quality and photorealism is highly comparable to the state-of-the-art.
-Furthermore, to further confirm the superiority of AccidentSim, we present qualitative results in Figure 3 to compare AccidentSim with baseline methods, and in Figure 4 to showcase generated videos across different road types.
-4.4 Post-Collision Trajectory Evaluation
-This experiment aims to confirm the physical plausibility and real-world consistency of post-collision vehicle trajectories generated by AccidentLLM, using quantitative comparisons against simulated ground truth.
-Using vehicle collision trajectories from CARLA as ground truth, we evaluate performance based on two criteria: Trajectory Prediction Error and Physical Consistency Error.
-Specifically, Trajectory Prediction Error is measured by the L2 distance between predicted and ground truth trajectories.
-Physical Consistency Error is assessed using the Mean Absolute Error (MAE) and Root Mean Square Error (RMSE) of impulse and momentum changes before and after the collision.
-The physical quantities of momentum and impulse are computed with reference to the models proposed by Zhou et al.
-[45].
-Given the limited prior work on vehicle collision trajectory prediction, we identify and adopt AutoVFX [17], which is a physics-based simulation method for generating collision videos, as the most suitable baseline due to its alignment with the physical modeling goals of our study.
-As shown in Table 3, AccidentSim consistently outperforms AutoVFX in both accuracy and physical consistency.
-It achieves lower average L2 errors across most scenarios, particularly in Straight Obstacle, Right Turn, and Lane Changing, demonstrating more stable and precise trajectory predictions.
-Moreover, AccidentSim shows significantly smaller errors in impulse and momentum metrics, indicating stronger adherence to physical laws such as conservation of momentum.
-Additional visualizations in Figure 3 demonstrate that the simulated vehicle collisions by AccidentSim are physically constrained by gravity and exhibit smoother trajectories more consistent with real-world dynamics.
-5 Conclusion
+To do this, we follow the setup from ChatScene [64] and fine-tune a model trained with Soft Actor-Critic from SafeBench [55], referred to as “Pre Fine-tuning” in Table IV, without AccidentSim’s generated collision videos for training.
+V Conclusion
 In this paper, we introduce AccidentSim, an innovative framework designed to generate vehicle collision videos with physically realistic collision trajectories.
@@
 Future work will integrate vehicle deformation and additional physical cues to improve realism.
-Supplementary Material
-6 Table of Contents
-  • Pre-Collision Trajectory Planning
-  • Information Extraction from Accident Reports
-  • Physical Simulation of Accident Scenario
-  • Additional Qualitative Analysis of Collision Dynamics
-  • Human and GPT-4o Evaluation
-  • Collision Reduction with AccidentSim
-Appendix A Algorithm for Pre-Collision Trajectory Planning
-The proposed algorithm for pre-collision trajectory planning, mentioned in Section 3.1 of the paper, is detailed in Algorithm 1.
-Figure 5: The process of extracting physical clues and contextual information from accident reports.
-Input : Vehicle motion information $\mathcal{I}$ (includes dynamics of all vehicles); scene map $\mathcal{M}$ containing lane nodes
-Output : Valid pre-collision trajectories $\mathcal{T}$; collision information $\mathcal{S}$
-Phase 1: Initial State Extraction
-  $\{\mathcal{C},n,D,V,A,\theta\}\leftarrow\textsc{ExtractState}(\mathcal{I})$ // Extract essential vehicle and collision parameters
-  $P_{0}\leftarrow\textsc{DetermineStartingPoints}(D,\theta)$ // Determine initial positions based on dynamics and directions
-Phase 2: Lane and Path Selection
-  $L\leftarrow\textsc{ParseMap}(\mathcal{M})$ // Parse scene map to identify candidate lanes
-  $\mathcal{L}\leftarrow\emptyset$ // Initialize set for valid lanes
-  foreach lane $l_{i}\in L$ do
-     if IsValidLane($l_{i}$, $\theta$, $\mathcal{C}$) then
-        $\mathcal{L}\leftarrow\mathcal{L}\cup\{l_{i}\}$ // Add valid lane $l_{i}$ to $\mathcal{L}$
- $\mathcal{G}\leftarrow\textsc{GatherLaneCombos}(\mathcal{L},n,D,V,A,\theta)$ // Form lane combinations for potential trajectories
-Phase 3: Trajectory Generation and Validation
-  foreach lane combination $G_{i}\in\mathcal{G}$ do
-     if MeetsCollisionCriteria($G_{i}$, $\mathcal{C}$) then
-        $G_{s}\leftarrow G_{i}$ // Select lane combination $G_{s}$ that satisfies collision criteria
-        break;
- $\mathcal{T}\leftarrow\textsc{GenerateTrajectory}(G_{s},P_{0})$ // Generate trajectories $\mathcal{T}$ based on selected lane combo $G_{s}$, and initial points $P_{0}$
- $\mathcal{S}\leftarrow\textsc{ImminentCollisionState}(\mathcal{T},V,A)$ // get imminent collision state $\mathcal{S}$ based on selected trajectories $\mathcal{T}$, and dynamics $\{V,A\}$
-return $\{\mathcal{T},\mathcal{S}\}$
-Algorithm 1 Pre-Collision Trajectory Planning
-Table 4: Information Extraction Accuracy: We compare AccidentSim with SoVAR[15], SoVAR_N [15], and AC3R[20] regarding their accuracy in extracting contextual information, such as road networks, traffic guidance, and dynamic object information from accident reports.
-Attributes Road Network and Traffic Guidance [15] Dynamic Objects [15]
-[…]
-SoVAR [15] 93.33% 100.00% 96.00% 76.33% 90.67% 100.00%
-SoVAR_N [15] 76.00% 91.33% 95.33% 55.00% 90.00% 100.00%
-AC3R [20] 70.67% 82.67% 78.67% 43.67% 11.67% 94.67%
-Ours 98.00% 100.00% 98.00% 93.00% 88.00% 100.00%
-Figure 6: Demonstration of physical simulations in CARLA for vehicle collisions.
-Figure 7: The collision videos generated by traditional methods such as ChatSim [34] exhibit model-interpenetration errors, where the two vehicles pass through each other.
-In contrast, AccidentSim, which incorporates physical constraints, produces physically realistic collision effects.
-Appendix B Information Extraction from Accident Reports
-The process of information extraction from the accident reports, mentioned in Section 3.2.1 of the paper, is detailed below.
-In AccidentSim, information related to the collision scene reconstruction is extracted from accident reports and analyzed using the Llama-3.1-8B model [11].
-Figure 5 shows an example of an accident report involving a collision between a Ford and a Dodge vehicle.
-In this example, AccidentSim extracts contextual information from the report, indicating that the collision occurred on a two-lane rural road with one right-curving lane in each direction and a speed limit of 89 km/h.
-The vehicles involved were a 2019 Ford F-450 Type-I ambulance and a Dodge Ram 2500 truck.
-Additionally, more detailed collision information is extracted from the report.
-For example, the report reveals that the left side of the Dodge vehicle collided with the left side of the Ford ambulance (collision type).
-At the moment of the collision, the vehicles were traveling at speeds of 49 km/h and 42 km/h, respectively.
-The collision occurred on a curve of the road (road type), with the Ford heading east and the Dodge heading west.
-The above extracted information enabled the reconstruction of road context and the replication of the collision within the CARLA simulator [10], as described in Section C of this Supplementary Material.
-We evaluated AccidentSim’s ability to extract collision information from accident reports against two mainstream methods: SoVAR [15] and AC3R [20].
-SoVAR [15] uses linguistic patterns to guide an LLM for extraction, while AC3R [20] uses NLP and hardcoded rules, requiring pre-segmented input for optimal performance.
-We also included the baseline SoVAR_N [15] in the evaluation, which directly extracts information from reports without using any linguistic patterns.
-As shown in the Table 4, AccidentSim achieves the highest accuracy for key attributes like lane number and speed limits, outperforming all the baselines.
-Furthermore, AccidentSim’s performance is also comparable to SoVAR for dynamic object information extraction.
-The superior extraction accuracy of AccidentSim is fundamental to ensure the accuracy of the subsequent simulation of vehicle collisions.
-Table 5: Comparison of Scenario Reconstruction Rate (SRR) between SoVAR and AccidentSim across different road types.
-Evaluation Metric Methods Road Types (Examples)
-Intersection T-junction Straight Road
-[…]
-AccidentSim 95.0% 76.3% 90.2%
-Appendix C Physical Simulation of Accident Scenario
-The physical simulation process for accident scenarios, mentioned in Section 3.2.2, is detailed as follows.
-As illustrated in Figure 6, we leverage key information extracted from accident reports, including physical parameters such as vehicle speed and position, as well as contextual information like road type and collision location, to conduct physical simulations of accidents within CARLA [10].
-Utilizing the extracted speed and position, we initialize vehicle states in CARLA and employ its automatic navigation capabilities to guide the vehicles toward the collision point.
-The simulation executes the collision event and captures precise data, such as the collision point and vehicle speed, via CARLA’s built-in collision sensors.
-This enables us to reproduce real-world accident scenarios with high physical fidelity, providing a reliable foundation for predicting novel collision trajectories using the fine-tuned AccidentLLM (as described in Section 3.2.3 of the paper).
-To quantitatively evaluate the effectiveness of AccidentSim in reconstructing accident scenarios, we adopted the methodology from SoVAR [15], using the Scenario Reconstruction Rate (SRR) as the primary evaluation metric.
-The SRR measures the overall proportion of trajectories generated from accident reports that successfully reproduce the reported collisions.
-Consequently, a higher SRR indicates superior accuracy in reconstructing real-world accident scenarios.
-As shown in Table 5, the results demonstrate that AccidentSim achieves higher reproduction accuracy than SoVAR across a range of randomly selected road configurations.
-Notably, the improvements are particularly significant for road types of T-intersections and straight roads.
-We attribute these advancements to AccidentSim’s enhanced trajectory planning capabilities resulting from the incorporation of physical constraints, which leads to greater accuracy and controllability in complex scenarios.
-Appendix D Additional Qualitative Analysis of Collision Dynamics
-Existing generative methods often struggle to produce physically realistic collision behaviors due to the lack of explicit physical constraints.
-To demonstrate and evaluate this, we compare AccidentSim with ChatSim [34], the first system that enables editable 3D driving scenes via natural language commands and external digital assets.
-As illustrated in Figure 7, ChatSim tends to exhibit severe errors of model interpenetration (i.e., vehicles passing through each other) during the collision.
-In contrast, AccidentSim leverages physical simulation to enforce rigid body dynamics and strict collision boundaries.
-The frame-by-frame analysis confirms that AccidentSim effectively eliminates such visual errors, yielding physically plausible outcomes that are consistent with real-world dynamics.
-Appendix E Human and GPT-4o Evaluation
-Figure 8: Prompt used for GPT-4o evaluation.
-Figure 9: Results of human evaluation.
-E.1 Human Evaluation
-The process of the human evaluation, mentioned in Section 4.3 of the paper, is detailed below.
-We recruited 30 participants aged 23 to 35 years, with a standard deviation of 3.70 years.
-The cohort comprised 20 males and 10 females, of whom 18 had at least two years of driving experience.
-Participants represented a range of professional backgrounds, including students, software professionals, financial sector employees, and transportation industry workers.
-We generated a total of 72 videos representing diverse vehicle collision scenarios, utilizing 6 methods as described in the paper, each covering 6 distinct collision types and 6 different scenarios.
-Each participant was requested to view and evaluate the generated collision videos.
-Participants first viewed each video and then completed a questionnaire, rating Physical Realism, Photorealism, and Semantic Consistency.
-To ensure the fairness and consistency of the evaluations, we provided detailed scoring criteria in the questionnaire, enabling participants to make informed and targeted assessments.
-The detailed evaluation results are presented in Figure 9.
-The experimental results show that AccidentSim achieves the best performance in terms of physical realism and semantic consistency, while being highly comparable to the state-of-the-art in photorealism.
-E.2 GPT-4o Evaluation
-The process of the GPT-4o evaluation, mentioned in Section 4.3 of the paper, is detailed below.
-We followed PhysGen3D [6] and employed GPT-4o as a multimodal evaluator to quantitatively assess the quality of the generated vehicle collision videos.
-This AI-based evaluation serves as an automated counterpart to human judgment, offering an additional layer of validation.
-Similar to the human evaluation, the evaluation centers on three core dimensions: Physical Realism, Photorealism, and Semantic Consistency.
-Each video is processed using a structured prompt specifically designed for GPT-4o.
-The prompt instructs the model to assign numerical scores from 1 to 5 to each of the three dimensions.
-The detailed definitions of these scores are aligned with the criteria used in the human evaluation to ensure consistency across assessment modalities.
-Figure 8 provides the prompt words we used when testing with GPT-4o.
-Appendix F Collision Reduction with AccidentSim
-For the experiment on collision reduction by AccidentSim, mentioned in Section 4.2 of the paper, the experimental configuration and training details are provided below.
-We conduct our experiments on the SafeBench platform [38] with an Nvidia A40 GPU.
-For scenario generation, we transfer vehicle driving situations from complex accidents reported by NHTSA to the driving scenarios in the Waymo Open Dataset and use the Scenic language [12] to script and simulate accident scenarios in Waymo environments within CARLA.
-In total, 4000 training scenarios and 1000 testing scenarios are generated.
-For the classification of accident scenarios, we adopt the accident scenario categorization proposed in ChatScene [41] and extend it to include three-vehicle and multi-vehicle collision scenarios.
-For training, we employ the SAC algorithm [16] to train the autonomous vehicle’s control policy.
-The SAC agent is trained for 500 epochs with a learning rate of 0.0001, and we evaluate its performance every 50 epochs to report the optimal results.